OpenStudy (anonymous):

Put the equation in the form ax^2+bx+c=0. (2x+1)^2-3x=5

OpenStudy (anonymous):

\[4x^2+4x+1-3x=5\] \[4x^2+x-4=0\]

OpenStudy (anonymous):

man you love these. \[(2x+1)^2-3x=5\] \[(2x+1)(2x+1)-3x=5\] \[4x^2+4x+1-3x=5\] \[4x^2+x+1=5\] \[4x^2+x-4=0\]

OpenStudy (anonymous):

NOOOO I hate these!